# SkillMimic: Learning Reusable Basketball Skills from Demonstrations

### v1

@@ 6.2. Metrics
6.3. Evaluating Skill Learning
In this section, we comprehensively evaluate the performance of SkillMimic in skill learning. We not only compare it with baseline methods but also conduct robustness tests on the skills acquired by SkillMimic. Additionally, we perform ablation studies and analyses from multiple dimensions, including reward function design and data volume. These experiments demonstrate the superior performance of SkillMimic, the rationality of its design, and its potential for further scaling up.
6.3.1. Comparing with Baselines.
Baseline Methods.
We compare our proposed SkillMimic with closely related imitation-based skill-learning methods, including DeepMimic (Peng et al., 2018) and AMP (Peng et al., 2021), which imitate body motion. We also create diverse versions of SkillMimic following the concept of reward designs in DeepMimic and AMP, dubbed SkillMimic-DM and SkillMimic-AMP, respectively. This allows for a more straightforward comparison of the differences in our reward function design. Detailed implementation of these methods can be found in the appendix.
Datasets.
We conduct experiments using data from three datasets: GRAB, BallPlay-V, and BallPlay-M. From the GRAB S8 subset, we select 5 cases, including grasping a cube, cylinder, flashlight, flute, and bottle. All data from BallPlay-V are used in the experiments. For BallPlay-M, we selected 40 pickup clips, one layup clip, one shot clip, and one dribble forward clip. These datasets were chosen for their representativeness and diversity, providing a comprehensive evaluation of our method’s overall performance.
Training Details.
Due to significant object and robot penetration issues in the GRAB and BallPlay-V data, using Reference State Initialization (RSI) often causes objects to bounce away. Therefore, for GRAB and BallPlay-V, we initialize the humanoid and object using the first frame of the reference clip and set the maximum simulation length to the clip length. For GRAB and BallPlay-V, we train 1 billion samples for all experiments, use the reference state as the condition, and only evaluate the imitation performance. For BallPlay-M, we apply RSI with a maximum simulation length of 60. We train approximately 0.65 billion samples for all experiments on BallPlay-M, each requiring around 10 hours on a single Nvidia RTX 4090 GPU. The exception is the pickup skill, which includes 40 data samples and is trained with about 3.2 billion samples.
Figure 7. A comparison of success rates across four typical basketball skills. Our method (SM) significantly outperforms skill-learning methods that only imitate body-only motions (DM and AMP). Additionally, our final reward design (SM) achieves superior overall performance compared to other variants (SM-DM and SM-AMP).
(a) 1 Clip, 0.5% Succ.
(b) 10 Clips, 12.2% Succ.
(c) 40 Clips, 76.5% Succ.
(d) 131 Clips, 85.6% Succ.
Figure 8. Pickup performance with different scales of training data. We collect 1, 10, 40, and 131 pickup clips from BallPlay-M to train pickup skills using SkillMimic. 1000 humanoids are initialized at the central, and 1000 balls are initialized at random locations 1 to 5 meters away from the center. The yellow dots indicate successfully picked balls and the green dots represent failures. Results demonstrate improved pickup skill performance with larger datasets.
Figure 9. Demonstration of the pickup skill learned from 40 HOI motion clips. Yellow denotes the initial frame. Left: The humanoid picks up a stationary ball effortlessly. Middle: The humanoid intercepts a ball with random velocity. Right: The humanoid adjusts after missing the ball initially (the frame in blue) and successfully retrieves it on the second attempt, showcasing the potential for learning robust and generalizable skills through extensive data collection.
Quantitative Results.
In Fig. 7, we present the success rate on 4 typical basketball skills from BallPlay-M. We can see that DeepMimic and AMP exhibit very low success rates in all skills. This is because they solely imitate body movements without object perception and constraints. When object perception and object reward functions are introduced, the re-implementation of DeepMimic-style and AMP-style rewards within our SkillMimic framework (SkillMimic-DM and SkillMimic-AMP) show significant performance improvements over their original versions, though the performance is still unstable. This instability is because they totally rely on kinematic imitation rewards, which do not adequately reflect the correctness of contact, leading to kinematic local optima. In contrast, our final reward design achieves stable performance across all skills. This stability is primarily because our reward function not only considers kinematics but also explicitly models contact and uses multiplicative methods to encourage balanced learning among various sub-rewards. Additionally, Tab. 1 presents detailed metrics of HOI imitation across two diverse datasets, GRAB and BallPlay-V, demonstrating that our method (SkillMimic) performs well even on datasets with kinematic biases. Comparatively, the absence of multiplicative rewards or CGR leads to noticeable performance drops (SkillMimic w/o multiplication and SkillMimic w/o CGR). This is because data with kinematic biases are more prone to kinematic local optima. Through these comprehensive evaluations, our method shows state-of-the-art performance in HOI imitation.
6.3.2. Evaluating Skill Robustness.
6.3.3. Ablation on Contact Graph Reward.
We conduct ablation experiments on diverse datasets to comprehensively evaluate the impact of Contact Graph Reward (CGR) on skill learning. Fig. 5 illustrates the qualitative results of ablating CGR on the GRAB and BallPlay-V datasets. It can be clearly observed that without CGR, the humanoid often resorts to incorrect contacts to achieve more stable object control, such as using its head to push the ball, clamping the ball with its hands and legs, or using its hands to support itself on a table for balance. These examples illustrate typical kinematic local optima. By incorporating CGR, these local optima are effectively eliminated, resulting in accurate interactions. Additionally, the variant SkillMimic w/o CGR in Tab. 1 presents the quantitative results of ablating CGR on two datasets. The results show that, without CGR, the contact error ($E_{\text{cg}}$) significantly increases, leading to kinematic local optima and causing a noticeable decline in overall imitation accuracy. These results underscore the critical importance of CGR for learning correct interactions.
6.3.4. Ablation on Data Volume.
To analyze the impact of data scale on the performance of skills acquired through SkillMimic, we conduct the following ablation experiments: (1) increasing the data for a single skill to evaluate the effect of data scale on its performance; (2) comparing the effectiveness of learning skills through independent training versus mixed training of multiple skills.
Data Scale Effect on a Single Skill.
We select the pickup skill for experiments due to its extensive semantic state transitions, as the ball can be in various conditions, leading to diverse state transitions for picking it up. We randomly select pickup clips from BallPlay-M to create four pickup skill collections: one with a single clip, one with 10 clips, one with 40 clips, and one with 131 clips. For each pickup skill collection, we train a single policy using SkillMimic for around 3.2 billion samples. During testing, balls are randomly distributed within a circular area with a radius of 1 to 5 meters centered around a humanoid. This setup is designed to evaluate the generalization performance of the learned pickup skills. Fig. 8 presents the results, where yellow dots indicate successfully picked-up balls and green dots represent failures. The results reveal a substantial improvement in pickup performance with increasing data scale. This improvement is attributed to the more comprehensive coverage of the state transitions inherent in the pickup skill provided by larger datasets. In Fig. 9, we also visualize how the pickup skill learned from 40 clips handles diverse ball states with diverse behavior. In the left graph, a ball is placed in a random stationary position, allowing a humanoid to run directly towards it and pick it up effortlessly. The middle graph introduces a scenario where the ball, starting from a random position, is given a random velocity. The humanoid dynamically alters direction to intercept and retrieve the moving ball. In the right graph, the ball’s velocity is increased, causing the humanoid to miss the ball on the first attempt. Then the humanoid quickly stands up, continues to pursue the ball, and successfully picks it up on the second attempt. These scenarios illustrate how extensive data collection enables the development of complex and adaptive ball retrieval skills. These experiments demonstrate the potential of SkillMimic to extend to learning more complex and diverse skills by enriching the dataset. Unlike previous approaches that required carefully designed skill-specific rewards, our purely data-driven method achieves a high success rate with diverse and adaptive behavior, which is particularly noteworthy.
(a) Manual Control
(b) Scoring
(c) Heading
(d) Circling
Figure 10. We train a single policy under a unified configuration to acquire various skills. These skills can be flexibly switched, as illustrated in (a), where yellow denotes shot, blue denotes pickup, and green denotes turnaround layup. Complex high-level tasks can be easily achieved by training a high-level controller (HLC) to manage skill switching: (b) scoring from random positions; (c) dribbling to target locations; and (d) dribbling with different radii.
Mixed Skill Training vs. Individual Skill Training.
We conduct the following experiment to analyze the effect between different skills when learned together. From BallPlay-M, we select one clip for the layup skill, one clip for the Dribble Left (DL) skill, one clip for the Dribble Right (DR) skill, and two clips for the Dribble Forward (DF) skill. These skills are commonly used and combined in real-world basketball games. We first train four individual policies for each skill independently, with each policy trained for around 0.65 billion samples. As a control, we then train these four skills using a single policy, and report the results when trained for around 0.65 billion samples (the same as the individual training) and 2.6 billion samples (4 times of the individual training, but the average sampling number for each skill is the same as in the individual training). During testing, we compare the success rates of executing each skill independently and transitioning between skills. Specifically, for testing each skill, we use the skill clips for Reference State Initialization (RSI) and execute the corresponding skill. For testing skill switching, we use the source skill clips for RSI and execute the target skill.
Success rates are calculated as described in Sec. 6.2. Tab. 3 presents the quantitative results. Despite equal sampling for each skill in both sets of experiments, mixed training shows a significant improvement in the success rates of individual skills and an even more substantial improvement in skill switching. It should be noted that while reward convergence is slightly faster in individual training, this approach is susceptible to overfitting. For example, the DL skill demonstrates excellent convergence of its reward, but due to the inadequate state cycle in the reference data, the DL skill trained independently tends to fall after a few dribbling steps, resulting in a zero success rate in sustained operations. Conversely, mixed training allows for cross-learning from other skills, thereby significantly enhancing the success rate of the DL skill. A similar phenomenon is observed in skill switching, where the reference data lacks examples of skill switches. Mixed training enables the policy to adapt to the state distributions of all skills, facilitating zero-shot skill switching during tests. These findings not only demonstrate that SkillMimic can support a single policy to learn various basketball skills but also underscore the importance of mixed training in enhancing skill generalization and robustness.
Table 3. Success rates of skills trained independently versus jointly. DF, DL, and DR denote Dribble Forward, Left, and Right. Ind. denotes individual training. 1$\times$ denotes 0.65 billion training samples while 4$\times$ denotes 4 times of that. Mixed training significantly improves both individual skill execution and skill switching, demonstrating the effectiveness of SkillMimic in handling diverse interaction skills at once.
Succ. on Individual Skills Succ. on Skill Switching
Training DF DL DR Layup DF-DL DF-DR DF-Layup DL-DF
[…]
6.4. Evaluating Complex Basketball Tasks
By training High-Level Controllers (HLC) to leverage skills learned via SkillMimic, humanoids can rapidly acquire complex basketball skills with simple task rewards. We conduct experiments on the four complex tasks described in Sec. 5.4. We employ a single skill policy as the skill prior, trained using SkillMimic for around 4.5 billion samples across seven skills: pickup, layup, turnaround layup, dribble left, dribble right, dribble forward, and shot. During training, the skill policy is kept fixed, and only the HLC is trained. The HLC operates at one-third the control frequency of the skill policy, i.e., 20 Hz. For tasks such as throwing, heading, and circling, we train for around 0.4 billion samples. For the more challenging scoring task, we train for around 1.2 billion samples. To objectively evaluate our method’s performance, we compared it against two baseline methods: (1) learning from scratch using PPO, and (2) learning with motion priors using ASE (Peng et al., 2022). The low-level controller of ASE is trained using the same data and training samples as our skill policy. For task training, all methods were trained using identical task rewards and simulation steps for fair comparisons.
Fig. 10 (a) showcases an example of manually controlling the pre-trained skill policy to perform skill switching. Specifically, we first execute the shooting skill, then switch to the pickup skill as the ball falls, and finally switch to the turnaround layup skill after retrieving the ball. Manual control, however, often fails to achieve precise goals, such as scoring a layup. Besides, since the success rate of skill switching is not 100%, effective skill switching requires switching within appropriate states, making manual control prone to potential failure. However, training a HLC to manage skill switching addresses these issues effectively.
Figure 11. Learning curves of different methods on 4 complex basketball tasks. On these challenging tasks, simple goal-related task rewards struggle to converge when training from scratch (PPO) or using motion priors (ASE), whereas our method achieves rapid convergence by using skill prior.
Table 4. Success rates on 4 complex basketball tasks. Both PPO (learn from scratch) and ASE (using motion prior) fail to converge on these four challenging tasks. In contrast, leveraging skill priors acquired through SkillMimic, our method effectively learns these difficult tasks.
PPO ASE Ours
[…]
Heading 0.70% 0.19% 93.04%
Circling 11.14% 4.37% 79.92%
Throwing 0.00% 0.00% 93.40%
Scoring 0.00% 0.00% 80.25%
Fig. 10 (b) shows several trajectories of the HLC performing skill controls to accurately deliver the ball to the target basket. Tab. 4 shows that our method achieves a layup success rate of up to 80.25%. Fig. 10 (c) presents a top-down view of the HLC controlling the humanoid to dribble the ball to target locations. Fig. 10 (d) depicts a top-down view of the HLC guiding the humanoid to dribble around a point with varying radii. As shown in Tab. 4, the success rates for the heading and circling tasks are 93.04% and 79.92%, respectively. These results highlight the extensive applicability of skill priors acquired through SkillMimic.
As shown in Fig. 11 and Tab. 4, both PPO and ASE exhibit poor performance in terms of both success rate and reward convergence. This is because the simple goal-defined task rewards are too difficult to learn for these complex basketball tasks. Like how human athletes, after mastering a wide array of basketball skills, can quickly adapt to new tasks, utilizing the pre-trained skill policy as a skill prior greatly simplifies the learning of complex tasks and results in rapid training convergence.
7. Discussion and Future Work
We presented a novel framework that learns various basketball skills in a purely data-driven manner. The proposed Contact Graph (CG) reward plays a critical role in human-object interaction (HOI) imitation, yielding significant improvements over traditional kinematic imitation rewards, particularly on data with kinematic errors. Thanks to a series of innovations, our approach is applicable to various basketball skills, enabling a single policy to master multiple skills and achieve zero-shot skill switching. Additionally, our method demonstrates a clear data scale effect, where the capability of the learned skills improves as the amount of data increases. We have also provided two basketball datasets to facilitate further research within the community. We believe this work paves an exciting new direction for humanoid skill learning and has the potential to extend to more general interactive skills.
Despite these merits, our method faces several limitations. Firstly, exploring more general skill learning within a single policy, such as household tasks or other types of sports, presents challenges in handling generalization across different objects, like varying object shapes. Moreover, perceiving multiple objects simultaneously requires a more general environmental perception approach, rather than directly obtaining privileged information for each object. Secondly, similar to previous works (Peng et al., 2021, 2022), our method predicts the next action based solely on the current moment without considering historical information. This can lead to convergence issues when learning from data with conflicting state transitions, such as a person standing for a long time before picking up a ball, where the data includes both standing-to-standing and standing-to-picking transitions, making it impossible for the policy to learn both transitions. In this work, we address this issue by dividing the conflicting data into different skills. In the future, we hope for more effective solutions, such as incorporating historical observations or learning automatic data segmentation and encoding.
Appendix A Hyperparameters
@@
Appendix B Details of Variant Models
B.1. Variants in Skill Learning
In the main paper, we developed multiple variants to enable thorough ablation studies and comparative analysis in skill learning. As a control measure, these variants are identical in all respects except for the variations in their reward functions. We will next delineate the specific differences in the reward functions of these variants. The symbols used for the sub-reward functions below are consistent with those in the main text. Unless specified, the hyperparameters of these sub-rewards are the same, as shown in Tab. 7.
SkillMimic-DM
@@ B.1. Variants in Skill Learning
(14) $r_{t}=r_{t}^{p}+r_{t}^{r}+r_{t}^{rv}+r_{t}^{o}.$
The hyperparameters of these sub-rewards are shown in Tab. 7.
SkillMimic-AMP
@@
(15) $r_{t}=-\text{log}\left(1-D(\boldsymbol{s}_{t},\boldsymbol{s}_{t+1})\right),$
where $\boldsymbol{s}$ represents the HOI state, $D$ denotes the discriminator.
@@
(17) $r_{t}=r_{t}^{b}*r_{t}^{o}*r_{t}^{rel}*r_{t}^{reg}.$
B.2. Variants in High-Level Tasks
To evaluate the performance of our method on high-level tasks, we established two sets of experiments for comparison: one involving training from scratch and the other utilizing ASE. For ASE, we trained a Low-Level Controller (LLC) using the same dataset as our skill policy, employing the following reward function:
(18) $r_{t}=-\text{log}\left(1-D(\boldsymbol{s}_{t},\boldsymbol{s}_{t+1})\right)+\beta\ \text{log}\ q\left(\boldsymbol{z}_{t}|\boldsymbol{s}_{t},\boldsymbol{s}_{t+1}\right).$
where $\boldsymbol{s}$ represents the body-only state (object is not considered), $D$ denotes the discriminator, $q$ denotes the encoder and $\boldsymbol{z}$ represents the latent code. The LLC can effectively generate body motions similar to those in the dataset.
Subsequently, we trained a High-Level Controller (HLC) using the same task reward as our method. The network architecture of this HLC is identical in size to that used in our approach; however, the HLC of ASE outputs continuous latent variables, whereas our HLC outputs discrete skill conditions.
Table 5. Hyperparameters for training skill policy.
@@ B.2. Variants in High-Level Tasks
Table 6. Hyperparameters for training high-level controller.
@@
Table 7. Hyperparameters of Sub-Rewards. SM denotes SkillMimic, and SM-DM denotes SkillMimic with DeepMimic-style rewards.
Parameter SM SM-DM

### v2

@@ 8.4 Success Rate
8.5 Hyperparameters
@@
8.6 Details of Compared Methods
8.6.1 Variants in Skill Learning
Since our approach is the first to learn interaction skills from demonstration, there are no direct benchmarks for comparison. Therefore, we adapt reward strategies commonly used in locomotion imitation, i.e., DeepMimic [36] and AMP [37], to an object-inclusive setting. Specifically, we adapt our reward to the styles of DeepMimic [36] and AMP [37] while keeping the other components unchanged for a fair comparison. We denote these variant versions as DeepMimic* and AMP*. We will next delineate the specific differences in the reward functions of these variants. The sub-rewards below are consistent to that defined in Sec. 8.2. Unless specified, the hyperparameters of these sub-rewards are the same, as shown in Tab. 9.
DeepMimic*
@@ 8.6.1 Variants in Skill Learning
$r_{t}=r_{t}^{p}+r_{t}^{r}+r_{t}^{rv}+r_{t}^{o}.$ (15)
The hyperparameters of these sub-rewards are shown in Tab. 9. Unlike DeepMimic [36], we do not incorporate phase information as policy input. This decision is made because phase-based methods face several practical limitations: they cannot operate continuously when reference data fails to form a complete cycle, and they show limited resistance to interference. All methods presented in this paper deliberately avoid using phase information.
AMP*
@@
$r_{t}=-\text{log}\left(1-D(\boldsymbol{s}_{t},\boldsymbol{s}_{t+1})\right),$ (16)
where $\boldsymbol{s}$ represents the HOI state which includes body and object state. $D$ denotes the discriminator.
@@
$r_{t}=r_{t}^{b}*r_{t}^{o}*r_{t}^{rel}*r_{t}^{reg}.$ (18)
@@
Table 7: Hyperparameters for training skill policy.
@@
Table 8: Hyperparameters for training high-level controller.
Parameter SM DM*
@@
Table 9: Hyperparameters of Sub-Rewards. SM denotes SkillMimic, and DM* denotes SkillMimic with DeepMimic-style rewards.
8.6.2 Variants in High-Level Tasks
To evaluate the performance of our method on high-level tasks, we established three sets of experiments for comparison: one involving training from scratch and the other two utilizing ASE [38], which first train a Low-Level Controller (LLC) using GAIL [19] then train a high-level controller to control the LLC. For fair comparison, we construct two versions of LLC for ASE. The first LLC follows the original style of ASE [38] which learns locomotion skills without considering object motion. We denote this method as ASE in our experiment. The second LLC follows the style of SkillMimic which learns interaction skills and considers object motion. We denote this object-inclusive LLC as ASE* in our experiment. The reward of both LLCs share the same formulation:
$r_{t}=-\text{log}\left(1-D(\boldsymbol{s}_{t},\boldsymbol{s}_{t+1})\right)+\beta\ \text{log}\ q\left(\boldsymbol{z}_{t}|\boldsymbol{s}_{t},\boldsymbol{s}_{t+1}\right),$ (19)
The difference is in the representations of $\boldsymbol{s}$. For ASE*, $\boldsymbol{s}$ contains the body and object state. For ASE, $\boldsymbol{s}$ represents the body-only state (object is not considered). $D$ denotes the discriminator, $q$ denotes the encoder and $\boldsymbol{z}$ represents the latent code.
Subsequently, we trained High-Level Controllers (HLC) to reuse these pre-trained LLCs for high-level tasks. For fair comparison, the task rewards are the same and the network size of the HLC is identical to that used in our approach. The HLC of ASE and ASE* outputs continuous latent variables, whereas our HLC outputs discrete skill conditions.